# itemizestditemize

### v2

@@ I Introduction
$LFC=\left\{d\in\mathbb{R}^{2}\;\bigg{|}\;\begin{array}[]{l}\nabla g_{i}^{\top}d\geq 0,i=1,2\\ \nabla g_{j}^{\top}d=0:j=3\end{array}\right\}.$ (6)
@@
After noticing the gap between the TC and the LFC, the reader might wonder why this is a problem. The reason lies in that, the equivalence of TC and LFC—typically ensured by CQs—is a prerequisite to establish the Karush–Kuhn–Tucker (KKT) conditions for local optimality [44]. Moreover, almost all primal-dual nonlinear programming (NLP) solvers (e.g., Snopt [25], Ipopt [54]) rely on the KKT optimality conditions—they search for a pair of primal and dual variables satisfying the KKT conditions. Therefore, the absence of CQs and the gap between TC and LFC breaks the theoretical foundations for primal-dual algorithms. The numerical consequence of this, as we will show in §IV, is that primal-dual solvers can exhibit poor convergence, stuck in infeasible solutions, and/or numerical instabilities when solving contact-implicit MPCC problems.
@@
$\min_{v,\lambda}\quad$ $J(v,\lambda),$ (8a)
@@
$0\leq\phi(x,\lambda)\perp\lambda\geq 0,$ (8e)
@@
where $v$ includes the trajectory of both robot state and control inputs, $\lambda$ denotes the complementarity variables typically associated with contact forces. We assume the objective function $J(v,\lambda)$ is convex—usually a quadratic loss function of tracking errors and control efforts. Constraint (8b) represents the nonlinear system dynamics. Constraints (8c) and (8d) are general inequality and equality constraints. The complementarity constraint (8e) enforces the relationship between $\lambda$ and the nonlinear function $\phi(v,\lambda)$. The expression “$\phi(x,\lambda)\perp\lambda$” should be interpreted as $\phi_{i}(x,\lambda)\cdot\lambda_{i}=0$, or its equivalent form $\phi_{i}(x,\lambda)\cdot\lambda_{i}\leq 0$ as suggested by [23], for every entry of $\phi(\cdot)$ and $\lambda$. It should be emphasized that problem (8) is nonconvex because we do not place any restrictions on the convexity of the constraints. We will give concrete examples of (8) in §IV. For now, let us work with the generic template (8).
@@
Our major contribution is to depart from the usual primal-dual algorithmic framework—due to the failure of CQs in MPCC—and propose a primal-only algorithm named CRISP (Robust Contact-Implicit motion planning with Sequential convex Programming). CRISP features a merit function with weighted $\ell_{1}$ penalty to measure primal feasibility and objective reduction. It solves a series of trust-region convex quadratic programs (QPs) formulated using local second-order information of the objective and first-order information of the constraints. In stark contrast with the well-known sequential quadratic programming (SQP) framework, CRISP (a) does not require second-order information of the constraints, and (b) does not maintain and update dual variables. Since the objective function is convex, the inner-loop QP subproblem inherits convexity by construction. This makes CRISP both simple to implement and computationally efficient.
@@
One then wonders whether a simple algorithm like CRISP can offer any convergence guarantees. We show that, somewhat surprisingly, the convexity of the objective function allows us to prove sufficient conditions under which CRISP can guarantee convergence to the stationary points of the merit function. Importantly, the sufficient conditions are numerically checkable, and hence the convergence can be “certified” from the numerical iterations of CRISP.
@@
Finally, we open-source a high-performance C++ implementation of CRISP. Our C++ implementation follows a generic nonlinear programming interface where the user defines objective function and constraints. We leverage automatic differentiation to obtain gradient and Hessian information. We then apply CRISP to solve five contact-implicit motion planning problems and benchmark its performance against both generic and robotics-specific solvers. We believe the way we model some of the contact-implicit planning problems is also new. Our numerical results demonstrate that CRISP consistently generates non-trivial and entirely new contact sequences from naive and even all-zero initializations, while existing solvers can struggle to find even feasible solutions.
@@
  • Benchmark. We model five contact-rich planning problems using the contact-implicit formulation and benchmark CRISP against existing solvers.
@@ II CRISP: Theory and Algorithm
Clearly, the contact-implicit planning problem (8) can be transformed into a generic nonlinear programming (NLP) problem, which also aligns with our implemented C++ solver. Let $x=[v,\lambda]$, we denote the NLP formulation of (8) as:
@@
Primal-only merit function. From a primal-only perspective, the core optimization challenge is to balance descent of the objective function with satisfaction of the constraints [44]. In CRISP, we adopt the following merit function with weighted $\ell_{1}$ penalty to evaluate the quality of a point $x$:
@@
We need to emphasize that, due to the lack of CQs in the MPCC problem, it is unclear whether minimization of the merit function (10) has the same type of guarantees as in the case where CQs hold. However, the intuition that the merit function balances objective reduction and constraint satisfaction still holds, and according to Example 2, the minimization of the merit function leads to the optimal solution, at least for the toy problem. We leave a precise investigation of the relationship between the merit function and the original problem in the case of MPCC as future work.
@@
From (II), the readers see that our method diverges from the subproblem formulation in the classical Sequential Quadratic Programming (SQP) methods [24, 21, 30], in the sense that (II) refrains from computing second-order derivatives of the constraints and the Lagrangian. Instead, we linearize the constraints while maintaining a quadratic model of the objective function only. Our approach aligns with the concept of sequential convex programming discussed in [15, 14, 41].
Leveraging convexity of the objective function $J$ and introducing the penalty term, the subproblem (II) is always convex and feasible. This offers several advantages. First, it reduces the computational burden associated with calculating second-order derivatives of constraints, which is particularly beneficial for problems with complex constraints (e.g., in robotics). Second, the convexity and guaranteed feasibility of the subproblem not only facilitate efficient solution of (II) using well-established convex optimization techniques but also enhance the robustness of the overall algorithm.
@@
where (12e) represents the $\ell_{\infty}$ norm constraint with trust-region radius $\Delta_{k}$. From (II) to (12), we have also introduced slack variables $v,w,t$ to reformulate the nonsmooth $\ell_{1}$ objective as smooth constraints. Evidently, the subproblem (12) is a standard convex quadratic program (QP).
@@
    The sequence of convex subproblems (12) converges, i.e., $x_{k}\rightarrow x^{*}$ for some point $x^{*}$
@@
Then, $x^{*}$ is a stationary point of the merit function (10).
@@ III CRISP: Implementation
User interface. CRISP follows the general nonlinear programming problem formulation in (9) where the user defines the objective function and constraints. Our implementation builds upon the fundamental data types in Eigen3 (vectors and matrices), wrapped with CPPAD [8] and CPPAD Code Generation (CG) [35]. This architecture allows users to define problems in an Eigen-style syntax, specifying objectives and constraints (both equality and inequality) while leveraging CPPAD for automatic derivative computation. The solver utilizes CPPAD to automatically obtain derivative information, while CG is employed to generate and store an auto-diff library. This library can be dynamically loaded for efficient evaluation of gradients and Hessians at specified points, significantly reducing computational overhead in subsequent solves. Besides, our implementation supports both parameterized and non-parameterized function definitions, allowing real-time modification of problem parameters (e.g., reference, terminal, or initial states) without the need to regenerate the library.
@@ IV CRISP: Benchmark
In this section, we present a comprehensive benchmark study to evaluate the performance of our proposed solver, CRISP, against several state-of-the-art optimization algorithms. As illustrated in Fig. 4, we have selected five representative contact-rich planning problems, each involving intricate interactions of normal (support) forces and friction forces. These problems showcase the diverse challenges encountered in practical applications where contact dynamics play a crucial role. All optimization problems are derived using the contact-implicit formulation and the detailed derivations are presented in Appendix B. The optimization horizon for all examples is 200 with time discretization $dt=0.02$ seconds. All experiments were conducted on a laptop equipped with a 13th Gen Intel(R) Core(TM) i9-13950HX processor.
@@ IV-A Benchmark Solvers
We compare CRISP against three state-of-the-art algorithms. The first is Snopt [25], a sparse nonlinear optimizer based on the SQP method. The second is Ipopt [54], an interior point optimizer designed for large-scale nonlinear programming. Finally, we include ProxNLP [29], s a primal-dual augmented Lagrangian algorithm designed by roboticists. As mentioned in the introduction, we exclude the bilevel contact-implicit algorithm [33] due to it only targeting at linear complementarity problems. Similarly, the ADMM-based contact-implicit solver in [6] is excluded for the same reason. Moreover, to the best of our knowledge, both [33] and [6] do not accept general nonlinear programming problem formulations.
@@ IV-B Benchmark Problems
    Push T: The modeling techniques for the push T problem are similar to those used in the push box problem. However, it presents its own challenges due to the non-convex shapes of the “T” and the increase in contact modes ($\lambda_{1}-\lambda_{8}$ as shown in Fig. 4(d)). This means we cannot simply construct complementarity constraints related to contact using four hyperplanes as we do with the push box problem. We now need to establish complementarity conditions that involve contact points lying on certain line segments and exerting corresponding contact forces, which requires special modeling techniques. Details of the problem formulation is demonstrated in Appendix B-D. Together with the slack variables, the total state dimension is 29.
@@
IV-C Numerical Results
@@
Performance metrics. We evaluate the solvers’ performance based on three key metrics. First, we examine the solution success rate, which reflects the ability of each solver to converge to a feasible solution from various initial guesses. Second, we assess the solution quality through the objective values achieved and the tracking error. Finally, we measure the solving efficiency by considering the number of iterations required and the total solving time. These metrics serve to answer a fundamental scientific question:
@@
Snopt $90$ $0.08$ $2.0\times 10^{-3}$ $4552.50$ $12.97$ $\bf 100$ $4.25$ $1.0\times 10^{-11}$ $3303.00$ $25.22$ $\bf{100}$ $1.7\times 10^{-4}$ $5.5\times 10^{-9}$ $5782.30$ $43.98$ $\bf{100}$ $0.67$ $\bf 6.4\times 10^{-8}$ $2257.50$ $210.92$
Ipopt $96.67$ $0.04$ $8.9\times 10^{-7}$ $\bf 288.21$ $\bf 0.51$ $\bf{100}$ $0.03$ $\bf 1.3\times 10^{-11}$ $\bf 407.69$ $0.92$ $\bf{100}$ $0.02$ $1.6\times 10^{-8}$ $404.96$ $1.36$ $\bf{100}$ $0.26$ $4.9\times 10^{-7}$ $1560.20$ $39.49$
PROXNLP $70$ $0.032$ $1.4\times 10^{-3}$ $716.25$ $140.92$ $66.67$ $0.03$ $3.6\times 10^{-5}$ $2000.00$ $1277.21$ $59.26$ $0.01$ $1.1\times 10^{-3}$ $361.92$ $1070.01$- -- -- -- --
@@
Quantitative results. For the first four examples, we recorded and summarized the numerical performance of the benchmark solvers under different initial states and initial guesses, as presented in the Table II and Fig. 5. We set a constraint violation threshold of $1\times 10^{-6}$ for each solver when running the experiments, and all final results with feasibility below $1\times 10^{-5}$ were considered successful, with their average constraint violation recorded. The objective function, as previously highlighted, primarily balances tracking error and control efforts. We analyze the performance of four solvers across three examples, focusing on median tracking error, average iteration count, and computation time. To better illustrate the performance and robustness of each solver under different initial guesses, we boxplot the distribution of constraint violations, objective values, and tracking errors in Fig. 5. These plots provide a comprehensive overview of solver behavior across all experiments, including both feasible and infeasible solutions, offering a complete picture of their performance characteristics.

### v3

@@ I Introduction
$LFC=\left\{d\in\mathbb{R}^{2}\;\bigg{|}\;\begin{array}[]{l}\nabla g_{i}^{\top}d\geq 0,i=1,2\\ \nabla g_{j}^{\top}d=0,j=3\end{array}\right\}.$ (6)
@@
After noticing the gap between the TC and the LFC, the reader might wonder why this is a problem. The reason lies in that, the equivalence of TC and LFC—typically ensured by CQs—is a prerequisite to establish the Karush–Kuhn–Tucker (KKT) conditions for local optimality [46]. Moreover, almost all primal-dual nonlinear programming (NLP) solvers (e.g., Snopt [24], Ipopt [57]) rely on the KKT optimality conditions—they search for a pair of primal and dual variables satisfying the KKT conditions. Therefore, the inherent absence of CQs and the gap between TC and LFC of the MPCC problems break the theoretical foundations for primal-dual algorithms. The numerical consequence of this, as we will show in §IV, is that primal-dual solvers can exhibit poor convergence, stuck in infeasible solutions, and/or numerical instabilities when solving contact-implicit MPCC problems.
@@
$\min_{v,\lambda}\quad$ $J(v,\lambda)$ (8a)
@@
$0\leq\phi(v,\lambda)\perp\lambda\geq 0,$ (8e)
@@
where $v$ includes the trajectory of both robot state and control inputs, $\lambda$ denotes the complementarity variables typically associated with contact forces. We assume the objective function $J(v,\lambda)$ is convex—usually a quadratic loss function of tracking errors and control efforts. Constraint (8b) represents the nonlinear system dynamics. Constraints (8c) and (8d) are general inequality and equality constraints. The nonlinear complementarity constraint (8e) enforces the relationship between $\lambda$ and the nonlinear function $\phi(v,\lambda)$. The expression “$\phi(x,\lambda)\perp\lambda$” should be interpreted as $\phi_{i}(x,\lambda)\cdot\lambda_{i}=0$, or its equivalent form $\phi_{i}(x,\lambda)\cdot\lambda_{i}\leq 0$ as suggested by [22], for every entry of $\phi(\cdot)$ and $\lambda$. It should be emphasized that problem (8) is nonlinear and nonconvex, as we impose no restrictions on the convexity of the constraints. This formulation is classified as a nonlinear complementarity problem (NCP) due to constraints (8e). A key distinction between NCP and linear complementarity problem (LCP) lies in the nonlinearity present in both the complementarity constraints and other constraints. While one can use $s=\phi(v,\lambda)$ to recast complementarity constraints as $\lambda\perp s$, this transformation merely shifts the nonlinearity to the additional constraint $s=\phi(v,\lambda)$. In the linear complementarity problem (LCP) literature, both $\phi$ and other constraints are typically required to be linear (as in [5] and LCQpow [28]). In contrast, CRISP is capable of handling generic nonlinear dynamics, as demonstrated in our examples. We will give concrete examples of (8) in §IV. For now, let us work with the generic template (8).
@@
Our major contribution is to depart from the usual primal-dual algorithmic framework—due to the failure of CQs in MPCC—and propose a primal-only algorithm named CRISP (Robust Contact-Implicit motion planning with Sequential convex Programming). CRISP features a merit function with weighted $\ell_{1}$ penalty to measure primal feasibility and objective reduction. It solves a series of trust-region convex quadratic programs (QPs) formulated using local second-order information of the objective and first-order information of the constraints. In stark contrast with the well-known sequential quadratic programming (SQP) framework, CRISP (a) does not require second-order information of the constraints, and (b) does not maintain and update dual variables. Since the objective function is convex, the inner-loop QP subproblem inherits convexity by construction. These make CRISP both simple to implement and computationally efficient.
@@
One then wonders whether a simple algorithm like CRISP can offer any convergence guarantees. We show that the convexity of the objective function allows us to prove sufficient conditions under which CRISP can guarantee convergence to the stationary points of the merit function. Importantly, the sufficient conditions are numerically verifiable, and hence the convergence can be “certified” from the numerical iterations of CRISP. Moreover, we show that the local minima of the merit function are indeed the local minimizers of the original problem if feasible for the original problem to close the loop.
@@
Finally, we open-source a high-performance C++ implementation of CRISP. Our C++ implementation follows a generic nonlinear programming interface where the user defines objective function and constraints. We leverage automatic differentiation to obtain gradient and Hessian information. We then apply CRISP to solve six contact-implicit motion planning problems and benchmark its performance against both generic and robotics-specific solvers. We believe the way we model some of the contact-implicit planning problems is also new. Our numerical results demonstrate that CRISP consistently generates non-trivial and entirely new contact sequences from naive and even all-zero initializations, while existing solvers can struggle to find even feasible solutions.
@@
  • Benchmark. We model six contact-rich planning problems using the contact-implicit formulation and benchmark CRISP against existing solvers.
@@ II CRISP: Theory and Algorithm
Clearly, the contact-implicit planning problem (8) can be transformed into a generic NLP problem, which also aligns with our implemented C++ solver. Let $x=[v,\lambda]$, we denote the NLP formulation of (8) as:
@@
Here, we absorb all other constraints into the general template $c$ for simplicity. Primal-only merit function. From a primal-only perspective, the core optimization challenge is to balance descent of the objective function with satisfaction of the constraints [46]. In CRISP, we adopt the following merit function with weighted $\ell_{1}$ penalty to evaluate the quality of a point $x$:
@@
We need to emphasize that, due to the lack of CQs in the MPCC problem, it is not fully clear whether minimization of the merit function (10) has the same type of guarantees as in the case where CQs hold. However, the intuition that the merit function balances objective reduction and constraint satisfaction still holds, and according to Example 2, the minimization of the merit function leads to the optimal solution, at least for the toy problem. We give a sufficient condition of local optimality in Proposition 9 and leave a precise investigation of the relationship between the merit function and the original problem in the case of MPCC for future work.
@@
From (II), the readers see that our method diverges from the subproblem formulation in the classical SQP methods [23, 20, 30], in the sense that (II) refrains from computing second-order derivatives of the constraints and the Lagrangian. Instead, we linearize the constraints while maintaining a quadratic model of the objective function only. Our approach aligns with the concept of sequential convex programming discussed in [14, 13, 43, 39, 38], including the well-known robotics package TrajOpt [51]. TrajOpt implements sequential convex programming for trajectory optimization and convex collision detection. However, it lacks user-friendly APIs to include contact constraints and does not support weighted $\ell_{1}$ merit functions or second-order corrections, nor does it provide convergence analysis. Beyond the SCP foundation, we aim for a high-performance implementation featuring carefully engineered modules and a generalized interface designed for broad contact-rich motion planning applications, while providing convergence guarantees in the contact-implicit context where CQs assumptions do not hold.
Leveraging the convexity of the objective function $J$ and introducing the penalty term, the subproblem (II) is always convex and feasible. This offers several advantages. First, it reduces the computational burden associated with calculating second-order derivatives of constraints, which is particularly beneficial for problems with complex constraints (e.g., in robotics). Second, the convexity and guaranteed feasibility of the subproblem not only facilitate efficient solution of (II) using well-established convex optimization techniques but also enhance the robustness of the overall algorithm.
@@
where (12e) represents the $\ell_{\infty}$ norm constraint with trust-region radius $\Delta_{k}$. From (II) to (12), we have also introduced slack variables $v,w,t$ to reformulate the nonsmooth $\ell_{1}$ objective as smooth constraints. Evidently, the subproblem (12) is a standard convex QP.
@@
    The sequence of convex subproblems (12) converges, i.e., $x_{k}\rightarrow x^{\star}$ for some point $x^{\star}$
@@
Then, $x^{\star}$ is a stationary point of the merit function (10).
@@
While the nonsmoothness of the merit function makes it challenging to verify whether its stationary points are local minima, a correspondence between the local minima of the merit function and the original problem can be established.
Proposition 9 (Local Optimality).
Let $x^{\star}$ be a local minimizer of the merit function $\phi_{1}(x;\mu)$ in (10). If $x^{\star}$ is feasible for the original problem (9), that is,
$\quad c_{i}(x^{\star})=0,\quad i\in\mathcal{E}$ (18)
$\quad c_{i}(x^{\star})\geq 0.\quad i\in\mathcal{I}$ (19)
Then $x^{\star}$ is also a local minimizer of the original problem (9).
Proof.
Since $x^{\star}$ is a local minimizer of $\phi_{1}(x;\mu)$, by definition, there exist $r>0~{}~{}s.t.~{}~{}\phi_{1}(x^{\star};\mu)\leq\phi_{1}(x;\mu),~{}~{}\forall||x-x^{\star}||<r$. Define the set $||x-x^{\star}||<r$ as $B_{r}(x^{\star})$ and let the feasible set of the original problem (9) be $\Omega$. Then, we have
$\phi_{1}(x^{\star};\mu)\leq\phi_{1}(x;\mu).~{}~{}\forall x\in B_{r}(x^{\star})\cap\Omega$
On the other hand, we have $\phi_{1}(x;\mu)=J(x)$ $\forall x\in\Omega$. Thus, evaluating the objective function of the original problem, we have:
$J(x^{\star})\leq J(x),~{}~{}\forall x\in B_{r}(x^{\star})\cap\Omega,~{}~{}x^{\star}\in\Omega$
which is the definition of a local minimizer. ∎
Combining Theorem 8 and Proposition 9, we conclude that if (a) CRISP converges to a stationary point $x^{\star}$ of the merit function, (b) $x^{\star}$ is a local minimizer of the merit function, and (c) $x^{\star}$ is feasible for the original problem (9), then $x^{\star}$ is also a local minimizer of (9). In other words, when CRISP computes a stationary point $x^{\star}$ of the merit function (which often occurs and can be certified) that is also locally optimal, the only way that $x^{\star}$ might fail to be a local minimizer of the original problem (9) is that $x^{\star}$ is infeasible for (9). We note that while this statement is theoretically useful, due to the nonsmoothness of the merit function, certifying local optimality of $x^{\star}$ for the merit function is challenging. We leave this for future work.
@@ III CRISP: Implementation
User interface. CRISP follows the general nonlinear programming problem formulation in (9) where the user defines the objective function and constraints. Our implementation builds upon the fundamental data types in Eigen3 (vectors and matrices), wrapped with CPPAD [7] and CPPAD Code Generation (CG) [35]. This architecture allows users to define problems in an Eigen-style syntax, specifying objectives and constraints (both equality and inequality) while leveraging CPPAD for automatic derivative computation. The solver utilizes CPPAD to automatically obtain derivative information, while CG is employed to generate and store an auto-diff library. This library can be dynamically loaded for efficient evaluation of gradients and Hessians at specified points, significantly reducing computational overhead in subsequent solves. Besides, our implementation supports both parameterized and non-parameterized function definitions, allowing real-time modification of problem parameters (e.g., reference, terminal, or initial states) without the need to regenerate the library. It is noteworthy to point that
@@ IV CRISP: Benchmark
In this section, we present a comprehensive benchmark study to evaluate the performance of our proposed solver, CRISP, against several state-of-the-art optimization algorithms. As illustrated in Fig. 4, we have selected six representative contact-rich planning problems, each involving intricate interactions of normal (support) forces and friction forces. These problems showcase the diverse challenges encountered in practical applications where contact dynamics play a crucial role. All optimization problems are derived using the contact-implicit formulation and the detailed derivations are presented in Appendix B. The optimization horizon for push T is 50 with time discretization $dt=0.05$ seconds, and all other examples are optimized over 200 steps with $dt=0.02$ seconds. All experiments were conducted on a laptop equipped with a 13th Gen Intel(R) Core(TM) i9-13950HX processor.
@@ IV-A Benchmark Solvers
In the first part (§IV-C), we compare CRISP against three state-of-the-art algorithms. The first is Snopt [24], a sparse nonlinear optimizer based on the SQP method. The second is Ipopt [57], an interior point optimizer designed for large-scale nonlinear programming. Finally, we include ProxNLP [29], a primal-dual augmented Lagrangian algorithm designed by roboticists. in subsection §IV-D, we present additional experiments to further demonstrate CRISP’s robustness and superior performance across a broader range of contact-rich motion planning scenarios. Specifically, (i) For the push T task, we reformulate the problem as a mixed-integer quadratically constrained problem (MIQCP), which we solve using Gurobi [27]. (ii) We evaluate the LCP solver LCQpow [28] on two examples with linear dynamics and linear complementarity constraints. (iii) We validate CRISP in real-world Push T with a Franka Panda robot arm.
@@ IV-B Benchmark Problems
    Push T: The modeling techniques for the push T problem are similar to those used in the push box problem. However, it presents its own challenges due to the non-convex shapes of the “T” and the increase in contact modes ($\lambda_{1}-\lambda_{8}$ as shown in Fig. 4(d)). This means we cannot simply construct complementarity constraints related to contact using four hyperplanes as we do with the push box problem. We now need to establish complementarity conditions that involve contact points lying on certain line segments and exerting corresponding contact forces, which requires special modeling techniques. Details of the problem formulation is demonstrated in Appendix B-D. Together with the slack variables, the total state dimension is 29. The experiment setting is the same with Push Box.
@@
IV-C Comparisons with Numerical Solvers
@@
Performance metrics. We evaluate the solvers’ performance based on three key metrics. First, we examine the solution success rate, which reflects the overall ability of each solver to converge to a feasible and informed solution from various settings (different initial/terminal conditions, different initial guesses). Second, we assess the quality of the solution through the objective values achieved and the tracking error. Finally, we measure the solving efficiency by considering the number of iterations required and the total solving time. These metrics serve to answer a fundamental scientific question:
@@
Snopt $63.33$ $0.08$ $2.0\times 10^{-3}$ $4552.50$ $12.97$ 00 $4.25$ $1.0\times 10^{-11}$ $3303.00$ $25.22$ $\bf{100}$ $1.7\times 10^{-4}$ $5.5\times 10^{-9}$ $5782.30$ $43.98$ $33.33$ $0.67$ $\bf 6.4\times 10^{-8}$ $2257.50$ $210.92$
Ipopt $86.67$ $0.04$ $8.9\times 10^{-7}$ $\bf 288.21$ $\bf 0.51$ $94.44$ $0.03$ $\bf 1.3\times 10^{-11}$ $\bf 407.69$ $0.92$ $87.50$ $0.02$ $1.6\times 10^{-8}$ $404.96$ $1.36$ $72.22$ $0.26$ $4.9\times 10^{-7}$ $1560.20$ $39.49$
PROXNLP $60$ $0.032$ $1.4\times 10^{-3}$ $716.25$ $140.92$ $66.67$ $0.03$ $3.6\times 10^{-5}$ $2000.00$ $1277.21$ $59.26$ $0.01$ $1.1\times 10^{-3}$ $361.92$ $1070.01$- -- -- -- --
@@
Quantitative results. For the first four examples, we recorded and summarized the numerical performance of the benchmark solvers under different initial states and initial guesses, as presented in the Table II and Fig. 5. Solutions with constraints violation below $10^{-5}$, translation error norm below 0.1, translation velocity error norm below 0.5, angular error below $\pi/6$, and angular velocity error below $0.1\pi\,\textup{rad}/\textup{s}$ as deemed successful. The objective function, as previously highlighted, primarily balances tracking error and control efforts. We analyze the performance of four solvers across the first four examples, focusing on median tracking error, average iteration count, and computation time. To better illustrate the performance and robustness of each solver under different initial guesses, we boxplot the distribution of constraint violations, objective values, and tracking errors in Fig. 5. These plots provide a comprehensive overview of solver behavior across all experiments, including both feasible and infeasible solutions, offering a complete picture of their performance characteristics.
@@
IV-D Additional Experiments
Additional Baselines on Push T. As an alternative formulation, we reformulated Push T as an MIQCP by associating each facet with a binary variable indicating the contact mode. Using Gurobi to solve this formulation, it failed to find solutions for planning horizons beyond 4 steps (within 10 minutes), making it incomparable with CRISP. The code to reproduce these experiments is open-sourced alongside CRISP.
LCP Solvers. The Transport and Waiter problems, characterized by linear dynamics and linear complementarity constraints, are naturally suited for a QP + LCP solver. We evaluated LCQpow [28] on these tasks. For the Transport problem, LCQpow achieves an average solve time of 251 s with an average tracking error of 0.4267, and 92.5% optimized trajectories are feasible. In comparison, as shown in Table II, CRISP significantly outperforms LCQpow with an average solve time of 0.77 s, an average tracking error of $1.1\times 10^{-4}$. For the more challenging Waiter problem, LCQPow fails to generate any feasible solutions. The code for reproducing these results is available in our open-source repository.